# ProveRAG: Provenance-Driven Vulnerability Analysis with Automated Retrieval-Augmented LLMs

### v1

@@ Abstract.
ProveRAG, LLM, Provenance, CVE, CWE, RAG, Vulnerability
copyright: nonedoi: .conference: ; ; $c$s: Information systems Language modelsccs: Information systems Top-k retrieval in databasesccs: Computing methodologies Natural language generationccs: Security and privacy Malware and its mitigation
1. Introduction
@@
In today’s fast-paced cybersecurity landscape, security analysts are continuously challenged to understand and mitigate newly discovered threats in real-time (Okutan et al., 2023). Common Vulnerabilities and Exposures (CVEs) (MITRE, 2024b) serve as a critical framework for identifying and detecting known vulnerabilities. With over 300,000 CVEs cataloged since 1999, the sheer volume makes it increasingly difficult and time-consuming for analysts to detect patterns and respond effectively to new and emerging threats, particularly when faced with unknown vulnerabilities. Public vulnerability databases, such as the NIST National Vulnerability Database (NVD) (of Standards and, NIST) and privately maintained repositories, predominantly rely on labor-intensive manual processes to analyze and categorize vulnerabilities. This reliance on human effort can lead to delays in identifying and addressing potential threats, making it difficult to keep pace with the rapidly evolving threat landscape, where timely analysis is critical for effective mitigation (Okutan et al., 2023). For example, the CVE-2024-0302 is a critical vulnerability that, if left unchecked, could lead to deserialization of untrusted data. To mitigate this vulnerability, fields should be made transient to be protected from deserialization. However, protecting an enterprise network necessitates discovering which vulnerabilities are relevant by combing through over 300,000 threats aggregated in these databases with more added each day. Separate mitigation databases, such as Common Weakness Enumeration (CWE) (MITRE, 2024a), must be explored to identify, implement, and deploy mitigations to the system. Due to the sheer volume of information, manually identifying and mitigating threats is challenging. In addition, automating the identification and mitigation process is essential for reducing the risk and enhancing overall resilience.
@@
Large Language Models (LLMs) have been widely adopted in cybersecurity applications (Mitra et al., 2024; Ullah et al., 2024; Cheshkov et al., 2023; Deng et al., 2024; Karlsen et al., 2024). They have been particularly effective in assisting with the analysis of Cyber Threat Intelligence (Mitra et al., 2024; Perrina et al., 2023), enhancing log-based anomaly detection (Karlsen et al., 2024; Qi et al., 2023), supporting LLM-assisted attacks (Sharma and Dash, 2023; Gupta et al., 2023; Deng et al., 2024), and improving vulnerability detection (Cheshkov et al., 2023; Ullah et al., 2024). However, prior work has identified three aspects that limit the broader adoption of LLMs to cybersecurity applications (Huang et al., 2023; Rawte et al., 2023; Zhu et al., 2024a; Liu et al., 2024), which we aim to address in this work. These limitations are: 1) Hallucination: the propensity to generate misleading information that appears plausible but is factually incorrect. Some works have shown the consequences of this problem (Huang et al., 2023; Rawte et al., 2023). 2) Temporal Knowledge: the failure of LLMs to respond accurately when they are exposed to data that comes after their training data’s cutoff (Zhu et al., 2024b, a). In cybersecurity, daily emerging vulnerabilities—over 25,000 identified in 2024 alone—pose serious risks, especially since they appeared after LLMs’ training data cutoff. Retrieval Augmented Generation (RAG) techniques have been proposed to address this problem (Borgeaud et al., 2022). However, the adoption of RAG for out-of-distribution data often leads to 3) Limited Context Windows: which restrict the number of tokens that can be provided to prompt the model, thereby limiting RAG-based temporal alignment approaches (Liu et al., 2024). These limitations constitute a major challenge of leveraging LLMs in cybersecurity, where accuracy, time, and up-to-date information is essential.
@@
In this work, we aim to improve the adaptation of LLMs in threat analysis. To do so, we propose ProveRAG, an advanced LLM-powered system designed to assist security teams in rapidly analyzing CVEs with verifiable evidence. ProveRAG goes beyond merely identifying vulnerabilities; it generates comprehensive exploitation steps and provides actionable mitigation strategies. What sets ProveRAG apart is its ability to cross-reference and validate information from authoritative sources like the NVD and CWE, ensuring that the insights it delivers are accurate and trustworthy. By doing so, we demonstrate that ProveRAG exhibits fewer instances of omission and hallucination through self-critiquing its own response. A key innovation in ProveRAG is its incorporation of provenance tracking into its analysis. By documenting the source of every piece of information, ProveRAG mitigates the risk of inaccuracies or hallucinations common in LLMs and gives security analysts the confidence to make swift, informed decisions.
@@
We test ProveRAG on 2024 CVEs, and we compare with two different baseline approaches: 1) prompt-only and 2) chunking retrieval during generation, and we compare it to our proposed summarizing technique (integrated into ProveRAG). The prompt-only technique is directly asking the LLM details about a specific CVE exploitation and mitigation strategies with data after its training cutoff knowledge. For the chunking technique, we enhance information retrieval by splitting the content into smaller chunks and embedding these chunks as vectors using a Transformer-based embedding model (OpenAI, 2024). These embeddings are then indexed to facilitate efficient retrieval. During runtime, the user’s query is also vectorized using the same encoder, and a similarity search is conducted against the indexed chunks. The top-k results are retrieved and fed as context into the LLM’s prompt, allowing the model to answer the user’s query with the most relevant content. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
Therefore, we design a summarizing technique to process CVEs, inspired by (Kim et al., 2024; Jiang et al., 2024). We start by exploring the NVD website and summarize the content with respect to a specific CVE-ID, then move to the CWE link, and relevant references (hyperlinks) to provide summary with respect to the exploitation and mitigation questions (emulating the typical workflow of a security analyst). A key advantage of the summarizing technique is its efficiency in handling large-scale data without relying extensively on document chunking, which enhances both retrieval quality and the accuracy of responses (while handling the context window limitation). This makes the system particularly suited for addressing more complex cybersecurity queries — such as those related to CVE mitigation that require integrating more external information.
@@
There is no easy way to determine whether a generated response is accurate or trustworthy. ProveRAG has the ability to provide verifiable evidences (TPs), detect hallucinations (FPs) and failed retrievals (FNs) for the LLM responses. This is called provenance where we carefully designed a structured, explicit, concise prompt with the chain-of-thoughts technique (Wei et al., 2022) to push the LLM into providing evidence for its responses. Note that we ensure that the LLM (gpt-4o-mini (OpenAI, 2024)) do not miss any essential information when finding evidence by utilizing almost the entire context window (128,000 tokens) for provenance. We further show that by providing the LLM with highly-structured context (Aqua Vulnerability Database (Security, 2024)), ProveRAG’s performance increases significantly. Our testing demonstrates that ProveRAG achieves outstanding detection accuracy, with 99% for summarization tasks and 97% for mitigation strategies. ProveRAG not only enables more reliable decision-making (compared to the chunking technique with over 30% increase in accuracy in providing mitigation strategies) but also meticulously documents its processes, ensuring a robust audit trail for future reviews. For evaluation, we ensure the coverage of every piece of information for each CVE when providing external resources (other than just NVD and CWE). This makes it an invaluable system for security analysts, empowering them to safeguard their systems with greater efficiency and accuracy. The contributions of our work is as follows:
@@
  • We demonstrate the effectiveness of ProveRAG for handling data that is out of the scope of LLMs’ training cutoff knowledge with a very high detection rate (i.e., TPs).
  • We indicate that ProveRAG is able to show proof for its responses (TPs), detect hallucinations (FPs), as well as demonstrating failed retrievals (FNs), which alleviates the hallucination and omission limitations of state-of-the-art LLMs.
@@
2. Background & Related Work
2.1. Large Language Models
@@
The significant advancements in Large Language Models is largely attributed to the integration of the Transformer architecture (Vaswani et al., 2017). These models are characterized by their extensive training on vast and diverse datasets, leading them to generate coherent and contextually relevant text that closely mirrors human language. OpenAI’s GPT-4o (OpenAI, 2024) and Meta’s LLama-3 (Dubey et al., 2024) are leading examples in this category. The ability of these models to manage diverse and complex tasks has been extensively documented in the literature (Min et al., 2023; Zhao et al., 2023). However, despite the efficacy of LLMs in diverse tasks, these models have shown to be prone to hallucination in their responses (Huang et al., 2023; Rawte et al., 2023).
2.2. Retrieval Augmented Generation
@@
The hallucination issue of LLMs stems from their reliance on pre-trained knowledge, which causes another limitation and that is the temporal scope of their training data. Such limitations pose significant challenges in fields requiring high precision, such as cybersecurity. To address these challenges and provide more factual knowledge, RAG was introduced (Borgeaud et al., 2022). RAG effectively combines search algorithms with LLM prompting: the model is prompted to answer a query using relevant information retrieved from external resources as context. This enhances LLMs’ ability to generate accurate and contextually relevant responses.
@@
Several advanced RAG techniques have been proposed to further refine this integration (Jiang et al., 2023; Su et al., 2024; Jeong et al., 2024; Asai et al., 2023; Kim et al., 2024; Jiang et al., 2024). FLARE (Jiang et al., 2023) improves model accuracy by predicting the next sentence and using low-confidence tokens as queries to re-retrieve relevant documents. DRAGIN (Su et al., 2024) leverages the LLM’s real-time information needs with self-attention to decide what and when to trigger retrieval. Adaptive-RAG (Jeong et al., 2024) employs a smaller LLM as a classifier to assess query complexity and subsequently selects the most appropriate retrieval strategy. Summarizing Retrievals (SuRe) (Kim et al., 2024) constructs summaries of retrieved passages for multiple answer candidates and confirms the most plausible answer by evaluating and ranking the generated summaries. TC-RAG (Jiang et al., 2024) integrates a memory stack system with backtracking and summarization functions to correct errors and eliminate redundant information. Despite these advancements, all these methods face the challenge of effectively showcasing the evidence underlying their generated responses.
2.3. LLMs in Vulnerability Assessment
@@
There are multiple works that integrated the use of LLMs in vulnerability assessment (Khare et al., 2023; Cheshkov et al., 2023; Du et al., 2024; Lu et al., 2024). Anton C. et al. (Cheshkov et al., 2023) used GPT models to identify CWE vulnerabilities in Java code and revealed that LLMs struggle with detecting vulnerable code. In contrast, Avishree K. et al. (Khare et al., 2023) demonstrated with multiple prompt strategies that LLMs can outperform deep learning models in vulnerability detection when carefully prompted, which indicates the importance of prompt engineering. Building on this, Lu et al. (Lu et al., 2024) introduced GRACE, a framework using in-context learning and graph-structured information with LLMs, which showed improvements in vulnerability detection on C/C++ datasets. Du et al. (Du et al., 2024) introduced Vul-RAG, LLM-based vulnerability detection technique that constructs a comprehensive knowledge base by extracting multi-dimensional insights from existing CVE instances. However, the reliance on human analysis in their approach makes it a manual time-consuming work and limits its efficiency in real-time vulnerability detection. In addition, PentestGPT (Deng et al., 2024) has shown capabilities in assisting in penetration testing. While it does not address the same problem this paper focuses on, the authors noted that LLMs are constrained by limited token sizes (i.e., context window), preventing long-term memory retention. This work builds on this observation and develop efficient RAG to bring LLM’s attention to relevant contents.
3. Motivation & Problem Formulation
@@
RQ2: How to manage information overload in the context of RAG-based LLMs for complex queries? Another significant problem we aim to solve is the issue of information overload (i.e., arising from context window limitations). This problem of inability of LLMs to process massive amounts of information and identify relevant information can lead to less effective responses, as discussed in literature (Liu et al., 2024). While more information is often seen as beneficial, our research shows that in the context of LLM-powered vulnerability analysis, an abundance of data can lead to inefficiencies (when chunking technique is used). The summarization technique we encountered in ProveRAG, can alleviate this issue. This technique can make the system especially effective in handling more complex cybersecurity issues, such as CVE mitigation queries that require more attention on additional relevant resources.
@@
RQ3: How to enhance the accuracy and trustworthiness of threat analysis while addressing issues related to hallucination and omission errors? One other key motivation for developing ProveRAG is the critical requirement to enhance the accuracy and trustworthiness of threat analysis by mitigating issues related to hallucination and omission errors. ProveRAG addresses this by integrating authoritative sources such as NVD and CWE into its framework. This integration ensures that the recommendations provided by ProveRAG are grounded in reliable and verifiable data, which we call provenance. To leverage the capabilities of LLMs, ProveRAG utilizes a structured, explicit, and concise prompt designed with a chain-of-thoughts technique (Wei et al., 2022). This approach compels the LLM to substantiate its responses with evidence, enhancing the credibility of the output. The system’s commitment to provenance is highlighted by its meticulous documentation of the sources of information it uses, creating a transparent and auditable trail. Based on our evaluation, this feature not only bolsters the analytical accuracy of ProveRAG but also fosters trust among security analysts by enabling them to verify the system’s recommendations.
Figure 1. ProveRAG: automated retrieval-augmented generation & self-critique provenance.
4. ProveRAG Methodology

### v2

@@ Abstract
1 Introduction
@@
In today’s fast-paced cybersecurity landscape, security analysts are continuously challenged to understand and mitigate newly discovered threats in real-time [24]. Common Vulnerabilities and Exposures (CVEs) [22] serve as a critical framework for identifying and detecting known vulnerabilities. With over 300,000 CVEs cataloged since 1999, the sheer volume makes it increasingly difficult and time-consuming for analysts to detect patterns and respond effectively to new and emerging threats, particularly when faced with unknown vulnerabilities. Public vulnerability databases, such as the NIST National Vulnerability Database (NVD) [23] and privately maintained repositories, predominantly rely on labor-intensive manual processes to analyze and categorize vulnerabilities. This reliance on human effort can lead to delays in identifying and addressing potential threats, making it difficult to keep pace with the rapidly evolving threat landscape, where timely analysis is critical for effective mitigation [24]. For example, the CVE-2024-0302 is a critical vulnerability that, if left unchecked, could lead to deserialization of untrusted data. To mitigate this vulnerability, fields should be made transient to be protected from deserialization. However, protecting an enterprise network necessitates discovering which vulnerabilities are relevant by combing through over 300,000 threats aggregated in these databases with more added each day. Separate mitigation databases, such as Common Weakness Enumeration (CWE) [21], must be explored to identify, implement, and deploy mitigations to the system. Due to the sheer volume of information, manually identifying and mitigating threats is challenging. In addition, automating the identification and mitigation process is essential for reducing the risk and enhancing overall resilience.
@@
Large Language Models (LLMs) have been widely adopted in cybersecurity applications [20, 33, 3, 4, 12]. They have been particularly effective in assisting with the analysis of Cyber Threat Intelligence [20, 26], enhancing log-based anomaly detection [12, 27], supporting LLM-assisted attacks [31, 7, 4], and improving vulnerability detection [3, 33]. However, prior work has identified three aspects that limit the broader adoption of LLMs to cybersecurity applications [8, 28, 37, 15], which we aim to address in this work. These limitations are: 1) Hallucination: the propensity to generate misleading information that appears plausible but is factually incorrect. Some works have shown the consequences of this problem [8, 28]. 2) Temporal Knowledge: the failure of LLMs to respond accurately when they are exposed to data that comes after their training data’s cutoff [38, 37]. In cybersecurity, daily emerging vulnerabilities—over 25,000 identified in 2024 alone—pose serious risks, especially since they appeared after LLMs’ training data cutoff. Retrieval Augmented Generation (RAG) techniques have been proposed to address this problem [2]. However, the adoption of RAG for out-of-distribution data often leads to 3) Limited Context Windows: which restrict the number of tokens that can be provided to prompt the model, thereby limiting RAG-based temporal alignment approaches [15]. These limitations constitute a major challenge of leveraging LLMs in cybersecurity, where accuracy, time, and up-to-date information is essential.
@@
In this work, we aim to improve the adaptation of LLMs in threat analysis. To do so, we propose ProveRAG, an advanced LLM-powered system designed to assist security teams in rapidly analyzing CVEs with verifiable evidence. ProveRAG goes beyond merely identifying vulnerabilities; it generates comprehensive exploitation steps and provides actionable mitigation strategies. What sets ProveRAG apart is its ability to cross-reference, rationalize, and validate information from authoritative sources like the NVD and CWE, ensuring that the insights it delivers are accurate and trustworthy. By doing so, we demonstrate that ProveRAG exhibits fewer instances of omission and hallucination through self-critiquing its own response. A key innovation in ProveRAG is its incorporation of provenance into its analysis. By documenting the source of every piece of information, ProveRAG mitigates the risk of inaccuracies or hallucinations common in LLMs and gives security analysts the confidence to make swift, informed decisions.
@@
We test ProveRAG on 2024 CVEs, and we compare with two different baseline approaches: 1) prompt-only and 2) chunking retrieval during generation, and we compare it to our proposed summarizing technique (integrated into ProveRAG). The prompt-only technique is directly asking the LLM details about a specific CVE exploitation and mitigation strategies with data after its training cutoff knowledge. For the chunking technique, we enhance information retrieval by splitting and embedding the content into smaller chunks. The top-k results are retrieved and fed as context into the LLM’s prompt, which allows the model to answer the user’s query with the most relevant content. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
Therefore, we design a summarizing technique to process CVEs, inspired by [14, 10]. We start by exploring the NVD website and summarize the content with respect to a specific CVE-ID. Next, we move to the associated CWE link, and further to the relevant references (hyperlinks) to provide summaries with respect to the exploitation and mitigation questions (emulating the typical workflow of a security analyst). A key advantage of the summarizing technique is its efficiency in handling large-scale data without relying extensively on document chunking, which enhances both retrieval quality and the accuracy of responses (while handling the context window limitation). This makes the system particularly suited for addressing more complex cybersecurity queries — such as those related to CVE mitigation that require integrating more external information.
@@
There is no easy way to determine whether a generated response is accurate or trustworthy. ProveRAG has the ability to provide verifiable evidences (TPs), detect hallucinations (FPs) and failed retrievals (FNs) for the LLM responses. This is called provenance where we carefully designed a structured, explicit, concise prompt with the chain-of-thought technique [35] to push the LLM into providing evidence for its responses. Note that we ensure that the LLM for ProveRAG (gpt-4o-mini [25]) do not miss any essential information when finding evidence by utilizing almost the entire context window (128,000 tokens) for provenance.
Moreover, we demonstrate that by providing the LLM with highly-structured context (Aqua Vulnerability Database [29]), ProveRAG’s performance increases significantly. Our testing demonstrates that ProveRAG achieves outstanding detection accuracy, with 99% for exploitation information and 97% for mitigation strategies. ProveRAG not only enables more reliable decision-making (compared to the chunking technique with over 30% increase in accuracy in providing mitigation strategies) but also meticulously documents its processes, ensuring a robust audit trail for future reviews. For evaluation, we ensure comprehensive coverage of all relevant information for each CVE by incorporating external resources beyond the NVD and CWE. Finally, we show that the proposed ProveRAG framework supports generalizability across different models by comparing OpenAI’s gpt-4o-mini and Meta’s Llama-3.1-8B [17]. This makes ProveRAG an invaluable system for security analysts, empowering them to safeguard their systems with greater efficiency and accuracy. The contributions of our work is as follows:
@@
  • We demonstrate the effectiveness of ProveRAG for handling data that is out of the scope of LLMs’ knowledge cutoff with a very high detection rate (i.e., TPs).
  • We show that ProveRAG is able to show proof for its responses (TPs), detect hallucinations (FPs), as well as demonstrating failed retrievals (FNs), which alleviates the hallucination and omission limitations of state-of-the-art LLMs.
  • We demonstrate that the ProveRAG framework enhances vulnerability analysis by incorporating an additional LLM (Llama-3.1-8B), thereby supporting the generalizability of the proposed approach.
@@
2 Background & Related Work
2.1 Large Language Models
@@
The significant advancements in Large Language Models is largely attributed to the integration of the Transformer architecture [34]. These models are characterized by their extensive training on vast and diverse datasets, leading them to generate coherent and contextually relevant text that closely mirrors human language. OpenAI’s GPT-4o [25] and Meta’s LLama-3 [6] are leading examples in this category. The ability of these models to manage diverse and complex tasks has been extensively documented in the literature [19, 36]. However, despite the efficacy of LLMs in diverse tasks, these models have shown to be prone to hallucination in their responses [8, 28].
2.2 Retrieval Augmented Generation
@@
The hallucination issue of LLMs stems from their reliance on pre-trained knowledge, which causes another limitation and that is the temporal scope of their training data. Such limitations pose significant challenges in fields requiring high precision, such as cybersecurity. To address these challenges and provide more factual knowledge, RAG was introduced [2]. RAG effectively combines search algorithms with LLM prompting: the model is prompted to answer a query using relevant information retrieved from external resources as context. This enhances LLMs’ ability to generate accurate and contextually relevant responses.
@@
Several advanced RAG techniques have been proposed to further refine this integration [11, 32, 9, 1, 14, 10]. FLARE [11] improves model accuracy by predicting the next sentence and using low-confidence tokens as queries to re-retrieve relevant documents. DRAGIN [32] leverages the LLM’s real-time information needs with self-attention to decide what and when to trigger retrieval. Adaptive-RAG [9] employs a smaller LLM as a classifier to assess query complexity and subsequently selects the most appropriate retrieval strategy. Self-RAG [1] employs self-reflection for retrieved documents to improve generation quality. However, Self-RAG trains two models–a critic model and a generator model– to assess its predictions. Our approach avoids any fine-tuning of LLMs, ensuring an automatic process designed to match the rapid pace in cybersecurity applications. Summarizing Retrievals (SuRe) [14] constructs summaries of retrieved passages for multiple answer candidates and confirms the most plausible answer by evaluating and ranking the generated summaries. TC-RAG [10] integrates a memory stack system with backtracking and summarization functions to correct errors and eliminate redundant information. Despite these advancements, all these methods face the challenge of effectively showcasing the evidence underlying their generated responses.
2.3 LLMs in Vulnerability Assessment
@@
There are multiple works that integrated the use of LLMs in vulnerability assessment [13, 3, 5, 16]. Anton C. et al. [3] used GPT models to identify CWE vulnerabilities in Java code and revealed that LLMs struggle with detecting vulnerable code. In contrast, Avishree K. et al. [13] demonstrated with multiple prompt strategies that LLMs can outperform deep learning models in vulnerability detection when carefully prompted, which indicates the importance of prompt engineering. Building on this, Lu et al. [16] introduced GRACE, a framework using in-context learning and graph-structured information with LLMs, which showed improvements in vulnerability detection on C/C++ datasets. Du et al. [5] introduced Vul-RAG, LLM-based vulnerability detection technique that constructs a comprehensive knowledge base by extracting multi-dimensional insights from existing CVE instances. However, the reliance on human analysis in their approach makes it a manual time-consuming work and limits its efficiency in real-time vulnerability detection. In addition, PentestGPT [4] has shown capabilities in assisting in penetration testing. While it does not address the same problem this paper focuses on, the authors noted that LLMs are constrained by limited token sizes (i.e., context window), preventing long-term memory retention. This work builds on this observation and develop efficient RAG to bring LLM’s attention to relevant contents.
3 Motivation & Problem Formulation
Figure 1: ProveRAG: automated retrieval-augmented generation & self-critique provenance.
@@ 3 Motivation & Problem Formulation
RQ2: How to manage information overload in the context of RAG-based LLMs for complex queries? Another significant problem we aim to solve is the issue of information overload (i.e., arising from context window limitations). This problem of inability of LLMs to process massive amounts of information and identify relevant information can lead to less effective responses, as discussed in literature [15]. While more information is often seen as beneficial, our research shows that in the context of LLM-powered vulnerability analysis, an abundance of data can lead to inefficiencies (when chunking technique is used). The summarization technique we encountered in ProveRAG, can alleviate this issue. This technique can make the system especially effective in handling more complex cybersecurity issues, such as CVE mitigation queries that require more attention on additional relevant resources.
@@
RQ3: How to enhance the accuracy and trustworthiness of threat analysis while addressing issues related to hallucination and omission errors? One other key motivation for developing ProveRAG is the critical requirement to enhance the accuracy and trustworthiness of threat analysis by mitigating issues related to hallucination and omission errors. ProveRAG addresses this by integrating authoritative sources such as NVD and CWE into its framework. This integration ensures that the recommendations provided by ProveRAG are grounded in reliable and verifiable data, which we call provenance. To leverage the capabilities of LLMs, ProveRAG utilizes a structured, explicit, and concise prompt designed with a chain-of-thought technique [35]. This approach compels the LLM to substantiate its responses with evidence, enhancing the credibility of the output. The system’s commitment to provenance is highlighted by its meticulous documentation of the sources of information it uses, creating a transparent and auditable trail. Based on our evaluation, this feature not only bolsters the analytical accuracy of ProveRAG but also fosters trust among security analysts by enabling them to verify the system’s recommendations.
4 ProveRAG Methodology
Figure 2: Chunking vs. Summarizing for retrieval process used in both generation and evaluation of ProveRAG.